OpenStudy (anonymous):

@zaynahf Could you please help me with this. I need the answer to this, please & thank you. :D Simplify the expression (-3)-4

OpenStudy (anonymous):

Is it \[-3^{-4}\] or -3 times -4?

OpenStudy (anonymous):

..And this is supposed to be in the Math section.. but since we already started, it's fine.

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

i am sorry ) :

OpenStudy (anonymous):

i just really need thee answer ) :

OpenStudy (anonymous):

@LyssaMeowPayne nah it's fine, i should have made sure, its my fault. When you have a negative exponent, do the reciprocal of the number. so if i had 2^-3, i would do: \[2^{-3}=\frac{ 1 }{ 2^{3} }=\frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

where did you get the two????

OpenStudy (anonymous):

??

OpenStudy (anonymous):

That was my example problem. Now use what i showed you to work out yours

OpenStudy (anonymous):

doesn't make sense

OpenStudy (anonymous):

Ok whats the base of your problem?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Ok, now whats the reciprocal of that?

OpenStudy (pratyush5):

(1/-3)^-4 = 1/ 81